# Phonon-mediated spin-polarized superconductivity in altermagnets

### v1

@@ Acknowledgements.
This method leaves, in principle, eleven free parameters, where $\eta$ is the bond strength between atoms of the same type, $\gamma$ is the NN strength, and $\rho$ is the NNN strength. To narrow down the number of free parameters, we consider the bond strength between all NN, etc., to be equal, hence $\eta_{i}^{j}=\eta$, $\gamma_{i}=\gamma$, and $\rho_{i}=\rho$. When determining the ratio of $\eta$, $\gamma$, and $\rho$, we assume the bond strength to decrease linearly with distance, such that $\gamma=2\eta$ and $\rho=\sqrt{2}\eta$. With these conditions, we only have one free parameter, $\eta$. We have chosen $\eta=-1.8\mathrm{N/m}$, so that the phonon bandwidth is about $30$ meV, see the main text. Also, due to the condition of equal bond strength for the three basis atoms, we have set all their masses equal, $M_{\alpha}=12~{}\mathrm{u}\,\forall\,\alpha$.
The rest of the FCM elements can be found by utilizing (S5), which gives
$\Phi^{11}\left(\left(0,2a\right)\right)=\sigma_{xy}^{\bm{\top}}\Phi^{33}\left(\left(2a,0\right)\right)\sigma_{xy},\quad$ $\Phi^{11}\left(\left(0,-2a\right)\right)=\sigma_{y}^{\bm{\top}}\Phi^{11}\left(\left(0,2a\right)\right)\sigma_{y},$
$\Phi^{11}\left(\left(-2a,0\right)\right)=\Phi^{11}\left(\left(2a,0\right)\right)^{\bm{\top}},\quad$ $\Phi^{22}\left(\left(0,2a\right)\right)=C_{4z+}^{\bm{\top}}\Phi^{22}\left(\left(2a,0\right)\right)C_{4z+},$
$\Phi^{22}\left(\left(-2a,0\right)\right)=C_{4z+}^{\bm{\top}}\Phi^{22}\left(\left(0,2a\right)\right)C_{4z+},\quad$ $\Phi^{22}\left(\left(0,-2a\right)\right)=C_{4z+}^{\bm{\top}}\Phi^{22}\left(\left(-2a,0\right)\right)C_{4z+},$
$\Phi^{33}\left(\left(-2a,0\right)\right)=\sigma_{x}^{\bm{\top}}\Phi^{33}\left(\left(2a,0\right)\right)\sigma_{x},\quad$ $\Phi^{33}\left(\left(0,2a\right)\right)=\sigma_{xy}^{\bm{\top}}\Phi^{11}\left(\left(2a,0\right)\right)\sigma_{xy},$
$\Phi^{33}\left(\left(0,-2a\right)\right)=\Phi^{33}\left(\left(0,2a\right)\right)^{\bm{\top}},\quad$ $\Phi^{12}\left(\left(2a,0\right)\right)=\sigma_{x}^{\bm{\top}}\Phi^{12}\left(\left(0,0\right)\right)\sigma_{x},$
$\Phi^{23}\left(\left(0,0\right)\right)=C_{4z+}^{\bm{\top}}\Phi^{12}\left(\left(0,0\right)\right)^{\bm{\top}}C_{4z+},\quad$ $\Phi^{23}\left(\left(0,-2a\right)\right)=C_{4z-}^{\bm{\top}}\Phi^{12}\left(\left(0,0\right)\right)^{\bm{\top}}C_{4z-},$
$\Phi^{13}\left(\left(2a,0\right)\right)=\sigma_{x}^{\bm{\top}}\Phi^{13}\left(\left(0,0\right)\right)\sigma_{x},\quad$ $\Phi^{13}\left(\left(0,-2a\right)\right)=\sigma_{y}^{\bm{\top}}\Phi^{13}\left(\left(0,0\right)\right)\sigma_{y}.$
$\Phi^{13}\left(\left(2a,-2a\right)\right)=\sigma_{\bar{xy}}^{\bm{\top}}\Phi^{13}\left(\left(0,0\right)\right)^{\bm{\top}}\sigma_{\bar{xy}}.$
@@
What remains to be determined are the three self-force elements, which are determined by using (S3), and are given by
$\Phi^{11}\left(\left(0,0\right)\right)$ $=-2\begin{pmatrix}\eta^{1}_{1}+\eta_{2}^{3}+\gamma_{1}+2\rho_{1}&0\\ 0&\eta^{1}_{2}+\eta^{3}_{1}+\gamma_{2}+2\rho_{1}\end{pmatrix},$
$\Phi^{22}\left(\left(0,0\right)\right)$ $=-2\begin{pmatrix}\eta^{2}_{1}+\eta_{2}^{2}+\gamma_{1}+\gamma_{2}&0\\ 0&\eta^{2}_{1}+\eta^{2}_{2}+\gamma_{1}+\gamma_{2}\end{pmatrix},$
$\Phi^{33}\left(\left(0,0\right)\right)$ $=-2\begin{pmatrix}\eta^{1}_{2}+\eta_{1}^{3}+\gamma_{2}+2\rho_{1}&0\\ 0&\eta^{1}_{1}+\eta^{3}_{2}+\gamma_{1}+2\rho_{1}\end{pmatrix}.$
[…]
$D(\textbf{k})=\begin{pmatrix}D^{11}(\textbf{k})&D^{12}(\textbf{k})&D^{13}(\textbf{k})\\ D^{12}(\textbf{k})^{\dagger}&D^{22}(\textbf{k})&D^{23}(\textbf{k})\\ D^{13}(\textbf{k})^{\dagger}&D^{23}(\textbf{k})^{\dagger}&D^{33}(\textbf{k})\end{pmatrix},$
$D^{11}_{xx}(\textbf{k})=2\left(\eta^{1}_{1}\left(\cos(2k_{x}a)-1\right)+\eta^{3}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-2\rho_{1}\right),$
$D^{11}_{yy}(\textbf{k})=2\left(\eta^{1}_{2}\left(\cos(2k_{x}a)-1\right)+\eta^{3}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{2}-2\rho_{1}\right),$
$D^{22}_{xx}(\textbf{k})=2\left(\eta^{2}_{1}\left(\cos(2k_{x}a)-1\right)+\eta^{2}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-\gamma_{2}\right),$
$D^{22}_{yy}(\textbf{k})=2\left(\eta^{2}_{2}\left(\cos(2k_{x}a)-1\right)+\eta^{2}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-\gamma_{2}\right),$
$D^{33}_{xx}(\textbf{k})=2\left(\eta^{3}_{1}\left(\cos(2k_{x}a)-1\right)+\eta^{1}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{2}-2\rho_{1}\right),$
$D^{33}_{yy}(\textbf{k})=2\left(\eta^{3}_{2}\left(\cos(2k_{x}a)-1\right)+\eta^{1}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-2\rho_{1}\right),$
@@
The dynamical matrix is Hermitian, so the elements that are not listed above and cannot be found by the constraint of hermicity are zero. The eigenvalue problem for the dynamical matrix is given by Eq. (1) in the main text.
S.X Electron-phonon coupling
@@
$\phi_{\textbf{r}_{i\alpha}}(\textbf{r})=\frac{1}{\pi^{3/4}\sqrt{\sigma_{x}\sigma_{y}\sigma_{z}}}\exp(-\frac{\left(x-r_{i\alpha}^{x}\right)^{2}}{2\sigma_{x}^{2}}-\frac{\left(y-r_{i\alpha}^{y}\right)^{2}}{2\sigma_{y}^{2}}-\frac{\left(z-r_{i\alpha}^{z}\right)^{2}}{2\sigma_{z}^{2}}),$ (S6)
@@
$t(\textbf{r}_{i\alpha}-\textbf{r}_{j\beta})=\int d^{3}r\phi_{\textbf{r}_{i\alpha}}(\textbf{r})A(\textbf{r})\phi_{\textbf{r}_{j\beta}}(\textbf{r})=A\exp(-\frac{{r_{ij\alpha\beta}^{x}}^{2}}{4\sigma_{x}^{2}}-\frac{{r_{ij\alpha\beta}^{y}}^{2}}{4\sigma_{y}^{2}}-\frac{{r_{ij\alpha\beta}^{z}}^{2}}{4\sigma_{z}^{2}}),$ (S7)
@@
$\bm{\nabla}_{r}t(\textbf{r}_{ij\alpha\beta})=-A\exp(-\frac{{r_{ij\alpha\beta}^{x}}^{2}}{4\sigma_{x}^{2}}-\frac{{r_{ij\alpha\beta}^{y}}^{2}}{4\sigma_{y}^{2}}-\frac{{r_{ij\alpha\beta}^{z}}^{2}}{4\sigma_{z}^{2}})\left(\frac{{r_{ij\alpha\beta}^{x}}^{2}}{2\sigma_{x}^{2}},\frac{{r_{ij\alpha\beta}^{y}}^{2}}{2\sigma_{y}^{2}},\frac{{r_{ij\alpha\beta}^{z}}^{2}}{4\sigma_{z}^{2}}\right),$ (S8)
where we note that at equilibrium, the z-component is zero. Such that to first order of the displacements, the electrons do not couple to out-of-plane phonon modes. We find the electron-phonon interaction by inserting this expansion in the electron Hamiltonian in the main text, Eq. (2). The first term in the expansion will simply give the electron hopping term, while the electron-phonon Hamiltonian is then given by
$H_{\mathrm{el-ph}}=\sum_{\langle i,j\rangle\sigma\alpha\beta}(\textbf{u}_{i\alpha}-\textbf{u}_{j\beta})\cdot\bm{\nabla}tc_{i\alpha\sigma}^{\dagger}c_{j\beta\sigma},$ (S9)
which, after inserting the Fourier transformed operators, quantifying the displacements, and utilizing the electron band operators as described in the main text, can be written as
$H_{\mathrm{el-ph}}=\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\sigma\lambda}^{(1)}(\textbf{k},\textbf{k+q})\left(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma},$ (S10)
@@
$g_{\sigma\lambda}^{(1)}(\textbf{k},\textbf{k+q})=\frac{t}{2}\sqrt{\frac{\hbar}{2N\omega_{\textbf{q}\lambda}}}\sum_{\alpha\beta}q_{\textbf{k}+\textbf{q},\sigma\alpha}^{*}q_{\textbf{k}\sigma\beta}\sum_{\bm{\delta}}e^{i\textbf{k}\cdot\bm{\delta}}\left(\frac{\delta_{x}}{\sigma_{x}^{2}},\frac{\delta_{y}}{\sigma_{y}^{2}}\right)\cdot\left[\frac{\bm{\hat{e}}_{\textbf{q}\beta}^{\lambda}}{\sqrt{M_{\beta}}}e^{i\textbf{q}\cdot\bm{\delta}}-\frac{\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}}{\sqrt{M_{\alpha}}}\right],$ (S11)
[…]

### v2

@@ Appendix B Phonon properties from symmetries
This method leaves, in principle, eleven free parameters, where $\eta$ is the bond strength between atoms of the same type, $\gamma$ is the NN strength, and $\rho$ is the NNN strength. To narrow down the number of free parameters, we consider the bond strength between all NN, etc., to be equal, hence $\eta_{i}^{j}=\eta$, $\gamma_{i}=\gamma$, and $\rho_{i}=\rho$. When determining the ratio of $\eta$, $\gamma$, and $\rho$, we assume the bond strength to decrease linearly with distance, such that $\gamma=2\eta$ and $\rho=\sqrt{2}\eta$. With these conditions, we only have one free parameter, $\eta$. We have chosen $\eta=-1.8\mathrm{N/m}$, so that the phonon bandwidth is about $30$ meV, as illustrated in Fig. 1(b). Also, due to the condition of equal bond strength for the three basis atoms, we have set all their masses equal, $M_{\alpha}=12~\mathrm{u}\,\forall\,\alpha$.
The rest of the FCM elements can be found by utilizing (15), which gives
$\Phi^{11}\left(\left(0,2a\right)\right)=\sigma_{xy}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{33}\left(\left(2a,0\right)\right)\sigma_{xy},$
$\Phi^{11}\left(\left(0,-2a\right)\right)=\sigma_{y}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{11}\left(\left(0,2a\right)\right)\sigma_{y},$
$\Phi^{11}\left(\left(-2a,0\right)\right)=\Phi^{11}\left(\left(2a,0\right)\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}},$
$\Phi^{22}\left(\left(0,2a\right)\right)=C_{4z+}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{22}\left(\left(2a,0\right)\right)C_{4z+},$
$\Phi^{22}\left(\left(-2a,0\right)\right)=C_{4z+}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{22}\left(\left(0,2a\right)\right)C_{4z+},$
$\Phi^{22}\left(\left(0,-2a\right)\right)=C_{4z+}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{22}\left(\left(-2a,0\right)\right)C_{4z+},$
$\Phi^{33}\left(\left(-2a,0\right)\right)=\sigma_{x}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{33}\left(\left(2a,0\right)\right)\sigma_{x},$
$\Phi^{33}\left(\left(0,2a\right)\right)=\sigma_{xy}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{11}\left(\left(2a,0\right)\right)\sigma_{xy},$
$\Phi^{33}\left(\left(0,-2a\right)\right)=\Phi^{33}\left(\left(0,2a\right)\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}},$
$\Phi^{12}\left(\left(2a,0\right)\right)=\sigma_{x}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{12}\left(\left(0,0\right)\right)\sigma_{x},$
$\Phi^{23}\left(\left(0,0\right)\right)=C_{4z+}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{12}\left(\left(0,0\right)\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}}C_{4z+},$
$\Phi^{23}\left(\left(0,-2a\right)\right)=C_{4z-}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{12}\left(\left(0,0\right)\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}}C_{4z-},$
$\Phi^{13}\left(\left(2a,0\right)\right)=\sigma_{x}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{13}\left(\left(0,0\right)\right)\sigma_{x},$
$\Phi^{13}\left(\left(0,-2a\right)\right)=\sigma_{y}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{13}\left(\left(0,0\right)\right)\sigma_{y}.$
$\Phi^{13}\left(\left(2a,-2a\right)\right)=\sigma_{\mkern 1.0mu\overline{\mkern-1.0muxy\mkern-1.0mu}\mkern 1.0mu}^{\raisebox{1.50694pt}{$\scriptstyle T$}}\Phi^{13}\left(\left(0,0\right)\right)^{\raisebox{1.50694pt}{$\scriptstyle T$}}\sigma_{\mkern 1.0mu\overline{\mkern-1.0muxy\mkern-1.0mu}\mkern 1.0mu}.$
@@
What remains to be determined are the three self-force elements, which are determined by using (13), and are given by
$\Phi^{11}_{\left(0,0\right)}$ $=-2\begin{pmatrix}\eta^{1}_{1}+\eta_{2}^{3}+\gamma_{1}+2\rho_{1}&0\\ 0&\eta^{1}_{2}+\eta^{3}_{1}+\gamma_{2}+2\rho_{1}\end{pmatrix}$,
$\Phi^{22}_{\left(0,0\right)}$ $=-2\begin{pmatrix}\eta^{2}_{1}+\eta_{2}^{2}+\gamma_{1}+\gamma_{2}&0\\ 0&\eta^{2}_{1}+\eta^{2}_{2}+\gamma_{1}+\gamma_{2}\end{pmatrix}$,
$\Phi^{33}_{\left(0,0\right)}$ $=-2\begin{pmatrix}\eta^{1}_{2}+\eta_{1}^{3}+\gamma_{2}+2\rho_{1}&0\\ 0&\eta^{1}_{1}+\eta^{3}_{2}+\gamma_{1}+2\rho_{1}\end{pmatrix}$.
This model can be extended to include bonds in the z-direction by adding the mirror symmetry of the z-plane, $\sigma_{z}$, in all $G_{I}$ sets, which will leave an additional independent parameter [78]. However, as we will see in Sec. C, the electron-phonon interaction with oscillations in the z-direction will be of higher order in the displacements, so we neglect these phonon modes. To solve for the phonon modes and frequencies, we turn to momentum-space, where the dynamical matrix is the Fourier transformation of the FCM, $D^{\alpha\beta}_{\mu\nu}(\textbf{k})\equiv\sum_{i}\Phi^{\alpha\beta}_{\mu\nu}(\textbf{R}_{i}-\textbf{R}_{j})e^{-i\textbf{k}\cdot(\textbf{R}_{i\alpha}-\textbf{R}_{j\beta})}$, which will then be a $6\crossproduct 6$ matrix with elements
$D(\textbf{k})$ $=\begin{pmatrix}D^{11}(\textbf{k})&D^{12}(\textbf{k})&D^{13}(\textbf{k})\\ D^{12}(\textbf{k})^{\raisebox{1.50694pt}{$\scriptstyle{\dagger}$}}&D^{22}(\textbf{k})&D^{23}(\textbf{k})\\ D^{13}(\textbf{k})^{\raisebox{1.50694pt}{$\scriptstyle{\dagger}$}}&D^{23}(\textbf{k})^{\raisebox{1.50694pt}{$\scriptstyle{\dagger}$}}&D^{33}(\textbf{k})\end{pmatrix}$,
$D^{11}_{xx}(\textbf{k})$ $=2\left[\eta^{1}_{1}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{3}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-2\rho_{1}\right],$
$D^{11}_{yy}(\textbf{k})$ $=2\left[\eta^{1}_{2}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{3}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{2}-2\rho_{1}\right],$
$D^{22}_{xx}(\textbf{k})$ $=2\left[\eta^{2}_{1}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{2}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-\gamma_{2}\right],$
$D^{22}_{yy}(\textbf{k})$ $=2\left[\eta^{2}_{2}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{2}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-\gamma_{2}\right],$
$D^{33}_{xx}(\textbf{k})$ $=2\left[\eta^{3}_{1}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{1}_{2}\left(\cos(2k_{y}a)-1\right)-\gamma_{2}-2\rho_{1}\right],$
$D^{33}_{yy}(\textbf{k})$ $=2\left[\eta^{3}_{2}\left(\cos(2k_{x}a)-1\right)\right$.
$\left.+\eta^{1}_{1}\left(\cos(2k_{y}a)-1\right)-\gamma_{1}-2\rho_{1}\right],$
@@
The dynamical matrix is Hermitian, so the elements that are not listed above and cannot be found by the constraint of hermicity are zero. The eigenvalue problem for the dynamical matrix is given by (1).
Appendix C Electron-phonon coupling
@@
$\phi_{\textbf{r}_{i\alpha}}(\textbf{r})=\frac{1}{\pi^{3/4}\sqrt{\sigma_{x}\sigma_{y}\sigma_{z}}}$ (16)
$\times\exp(-\frac{\left(x-r_{i\alpha}^{x}\right)^{2}}{2\sigma_{x}^{2}}-\frac{\left(y-r_{i\alpha}^{y}\right)^{2}}{2\sigma_{y}^{2}}-\frac{\left(z-r_{i\alpha}^{z}\right)^{2}}{2\sigma_{z}^{2}}),$
@@
$t(\textbf{r}_{i\alpha}-\textbf{r}_{j\beta})=\int d^{3}r\phi_{\textbf{r}_{i\alpha}}(\textbf{r})A(\textbf{r})\phi_{\textbf{r}_{j\beta}}(\textbf{r})$ (17)
$=A\exp(-\frac{{r_{ij\alpha\beta}^{x}}^{2}}{4\sigma_{x}^{2}}-\frac{{r_{ij\alpha\beta}^{y}}^{2}}{4\sigma_{y}^{2}}-\frac{{r_{ij\alpha\beta}^{z}}^{2}}{4\sigma_{z}^{2}})$ ,
@@
$\bm{\nabla}_{r}t(\textbf{r}_{ij\alpha\beta})=-A$ $\exp(-\frac{{r_{ij\alpha\beta}^{x}}^{2}}{4\sigma_{x}^{2}}-\frac{{r_{ij\alpha\beta}^{y}}^{2}}{4\sigma_{y}^{2}}-\frac{{r_{ij\alpha\beta}^{z}}^{2}}{4\sigma_{z}^{2}})$
$\hskip 5.59721pt\times\left(\frac{{r_{ij\alpha\beta}^{x}}}{2\sigma_{x}^{2}},\frac{{r_{ij\alpha\beta}^{y}}}{2\sigma_{y}^{2}},\frac{{r_{ij\alpha\beta}^{z}}}{4\sigma_{z}^{2}}\right),$ (18)
where we note that at equilibrium, the z-component is zero. Such that to first order of the displacements, the electrons do not couple to the OOP phonon modes. We find the electron-phonon interaction by inserting this expansion in the electron Hamiltonian in (2). The first term in the expansion will simply give the electron hopping term, while the electron-phonon Hamiltonian is then given by
$H_{\mathrm{el-ph}}=\sum_{\langle i,j\rangle\sigma\alpha\beta}(\textbf{u}_{i\alpha}-\textbf{u}_{j\beta})\cdot\bm{\nabla}tc_{i\alpha\sigma}^{\dagger}c_{j\beta\sigma},$ (19)
which, after inserting the Fourier transformed operators, quantifying the displacements, and utilizing the electron band operators as described in Sec. IV, can be written as
$H_{\mathrm{el-ph}}=\sum_{\textbf{k}\textbf{q}\sigma\lambda}g_{\sigma\lambda}^{(1)}(\textbf{k},\textbf{k}+\textbf{q})\left(a_{-\textbf{q},\lambda}^{\dagger}+a_{\textbf{q}\lambda}\right)d_{\textbf{k}+\textbf{q},\sigma}^{\dagger}d_{\textbf{k}\sigma},$ (20)
@@
$g_{\sigma\lambda}^{(1)}(\textbf{k},\textbf{k}+\textbf{q})=\frac{t}{2}\sqrt{\frac{\hbar}{2N\omega_{\textbf{q}\lambda}}}\sum_{\alpha\beta}q_{\textbf{k}+\textbf{q},\sigma\alpha}^{*}q_{\textbf{k}\sigma\beta}$ (21)
$\times\sum_{\bm{\delta}}e^{i\textbf{k}\cdot\bm{\delta}}\left(\frac{\delta_{x}}{\sigma_{x}^{2}},\frac{\delta_{y}}{\sigma_{y}^{2}}\right)\cdot\left[\frac{\bm{\hat{e}}_{\textbf{q}\beta}^{\lambda}}{\sqrt{M_{\beta}}}e^{i\textbf{q}\cdot\bm{\delta}}-\frac{\bm{\hat{e}}_{\textbf{q}\alpha}^{\lambda}}{\sqrt{M_{\alpha}}}\right]$ ,
where $\bm{\delta}$ are the equilibrium NN vectors. The same procedure can be done for the NNN hopping $t_{2}$, which gives an electron-phonon coupling $g_{\sigma\lambda}^{(2)}(\textbf{k},\textbf{k}+\textbf{q})$, that is almost identical to (21), only differing with the substitutions $t\rightarrow t_{2}$ and $\bm{\delta}\rightarrow\bm{\delta}_{2}$, where $\bm{\delta}_{2}$ are the vectors between NNNs at equilibrium. The total electron-phonon coupling is then given by $g_{\sigma\lambda}(\textbf{k},\textbf{k}+\textbf{q})\equiv g_{\sigma\lambda}^{(1)}(\textbf{k},\textbf{k}+\textbf{q})+g_{\sigma\lambda}^{(2)}(\textbf{k},\textbf{k}+\textbf{q})$.
C.1 Effective electron-electron interaction